\begin{thebibliography}{99.}%
% and use \bibitem to create references.
%
% Use the following syntax and markup for your references if 
% the subject of your book is from the field 
% "Mathematics, Physics, Statistics, Computer Science"
%
% Contribution 
\bibitem{review}
S.~Nojiri and S.~D.~Odintsov, %``Introduction to modified gravity and gravitational alternative for dark energy,''
 % eConf {\bf C0602061}, 06 (2006)
  Int.\ J.\ Geom.\ Meth.\ Mod.\ Phys.\  {\bf 4}, 115 (2007);   %[arXiv:hep-th/0601213]; arXiv:0807.0685 [hep-th]; arXiv:1011.0544 [hep-th];\\
  %S.~'i.~Nojiri and S.~D.~Odintsov,
  %``Unified cosmic history in modified gravity: from F(R) theory to Lorentz non-invariant models,''
 %  Phys.\ Rept.\  {\bf 505}, 59 (2011); % [arXiv:1011.0544 [gr-qc]].
T.~P.~Sotiriou and V.~Faraoni, Rev. Mod. Phys. \textbf{82} 451 (2010);
  %``f(R) Theories Of Gravity,'' %  [arXiv:0805.1726 [gr-qc]]; \\
F.~S.~N.~Lobo,   %``The dark side of gravity: Modified theories of gravity,''
  arXiv:0807.1640 [gr-qc]; %\\ 
%   T.~Harko and F.~S.~N.~Lobo,
  %``Geodesic deviation, Raychaudhuri equation, and tidal forces in modified gravity with an arbitrary curvature-matter coupling,''
 % arXiv:1210.8044 [gr-qc].
 %%CITATION = ARXIV:1210.8044;%%
%
%Extended Theories of Gravity.
% S.  Capozziello, M.  De Laurentis, Physics Reports \textbf{ 509} 167-321  (2011). %[arXiv:1108.6266 [gr-qc]] .
%
%\bibitem{book}
  S. Capozziello  and  V. Faraoni, {\it Beyond Einstein Gravity}, %Fundamental Theories of Physics Vol. 170, 
  Springer Ed., Dordrecht  (2011).
 %
%\bibitem{Ref1} 
A. Dobado and  A. L. Maroto {\it Phys. Rev.} {\bf D 52} 1895 (1995).
%
%\bibitem{Ref2} 
G. Dvali, G. Gabadadze and M. Porrati {\it Phys. Lett.} {\bf B 485} 208 (2000).
%
%\bibitem{Ref3} 
J. Beltr\'an and A. L. Maroto {\it Phys. Rev.}  {\bf D 78} 063005 (2008); {\it JCAP} {\bf 0903} 016 (2009)
; {\it Phys. Rev.} {\bf D 80 } 063512 (2009);  {\it Int. J. Mod. Phys.} {\bf D 18}  2243-2248 (2009).

\bibitem{unification}
  S.~Nojiri and S.~D.~Odintsov,
  %``Modified f(R) gravity consistent with realistic cosmology: From matter dominated epoch to dark energy universe,''
  Phys.\ Rev.\ D\ {\bf 74} 086005 (2006); %  [hep-th/0608008];
%Gauss-Bonnet gravity
E.~Elizalde and D.~S\'aez-G\'omez,
  %``F(R) cosmology in presence of a phantom fluid and its scalar-tensor counterpart: Towards a unified precision model of the universe evolution,''
  Phys.\ Rev.\ D\ {\bf 80}, 044030  (2009); 
  %[arXiv:0903.2732 [hep-th]].
S.~Nojiri, S.~D.~Odintsov and D.~S\'aez-G\'omez,
  %``Cosmological reconstruction of realistic modified F(R) gravities,''
  Phys.\ Lett.\ B\ {\bf 681}, 74  (2009); %   [arXiv:0908.1269 [hep-th]].
 % 
%\cite{Appleby:2009uf} \bibitem{Appleby:2009uf} 
  S.~A.~Appleby, R.~A.~Battye and A.~A.~Starobinsky,
  %``Curing singularities in cosmological evolution of F(R) gravity,''
  JCAP {\bf 1006}, 005 (2010); %  [arXiv:0909.1737 [astro-ph.CO]].
  %
%    A.~Dev, D.~Jain, S.~Jhingan, S.~Nojiri, M.~Sami and I.~Thongkool,
  %``Delicate f(R) gravity models with disappearing cosmological constant and observational constraints on the model parameters,''
  % Phys.\ Rev.\ D {\bf 78}, 083515 (2008).
%  [arXiv:0807.3445 [hep-th]].
  %%CITATION = ARXIV:0807.3445;%%

\bibitem{Perturbations}  
J. M. Bardeen, Phys. Rev. D22 1882 (1980);
S. M. Carroll, I. Sawicki, A. Silvestri and M. Trodden, New J. Phys. 8 323, (2006); 
Y. S. Song, W. Hu and I. Sawicki, Phys. Rev. D 75 044004, (2007);
A. A. Starobinsky, JETP Lett. 86, 157 (2007);
% R. Bean {\it et al.}, %D. Bernat, L. Pogosian, A. Silvestri and M. Trodden. 
% Phys. Rev. D 75 064020, (2007);
 S. Carloni, P. K. S. Dunsby and A. Troisi, Phys. Rev. D 77 024024, (2008); 
% S. Tsujikawa, Phys. Rev. D77:023507, (2008);
% S. Tsujikawa, K. Uddin, R. Tavakol, Phys. Rev. D 77 043007, (2008); 
A. Abebe, M. Abdelwahab, A. de la Cruz-Dombriz and P. K. S. Dunsby,  Class.  Quant. Grav.  {\bf 29} 135011 (2012). %XXXXX   {it et al.}, arXiv:1110.1191 [gr-qc];

\bibitem{PRD_Dombriz_2008} A. de la Cruz-Dombriz, A. Dobado and A. L. Maroto, Phys. Rev. D 77 123515 (2008).

\bibitem{PRL_Dombriz_2009} A. de la Cruz-Dombriz, A. Dobado and A. L. Maroto,  Phys. Rev. Lett. D 103, 179001 (2009).
  

 \bibitem{Dombriz-Saez}
 A.~de la Cruz-Dombriz and D.~S\'aez-G\'omez,
  %``On the stability of the cosmological solutions in $f(R,G)$ gravity,''
Class. Quant. Grav., {\bf 29} 245014, 2012.
% arXiv:1112.4481 [gr-qc].  
 
\bibitem{BH}
G. J. Olmo, Phys. Rev. D 75, 023511 (2007);
A. M. Nzioki, S. Carloni, R. Goswami and P. K. S. Dunsby, Phys. Rev. D81 084028 (2010);
A. de la Cruz-Dombriz, A. Dobado and A. L. Maroto, Phys. Rev. D 80, 124011 (2009) [Erratum: Phys. Rev. D 83, 029903(E) (2011)];  J.\ Phys.\ Conf.\ Ser.\  {\bf 229} (2010) 012033;
J. A. R. Cembranos, A. de la Cruz-Dombriz and  P. Jimeno Romero, e-Print: arXiv:1109.4519 [gr-qc]; AIP Conf. Proc. 1458, 439 (2012).

\bibitem{CMB-ERE2011} H. Bourhrous, A. de la Cruz-Dombriz, and P. K. S. Dunsby, AIP Conf. Proc. {\bf 1458}, 343 (2012).

 \bibitem{PRD-CMB-Dombriz} M. Abdelwahab, H. Bourhrous, A. de la Cruz-Dombriz, and P. K. S. Dunsby, in preparation.

\bibitem{Strings}
%   \bibitem{Bevis:2006mj} 
  N.~Bevis, M.~Hindmarsh, M.~Kunz and J.~Urrestilla,   %``CMB power spectrum contribution from cosmic strings using field-evolution simulations of the Abelian Higgs model,''
  Phys.\ Rev.\ D {\bf 75}, 065015 (2007); % [astro-ph/0605018].
% 
% \bibitem{Bevis_predictions} %Predicted Constraints on Cosmic Strings CMB
  S.~Foreman, A.~Moss and D.~Scott,
  %``Predicted Constraints on Cosmic String Tension from Planck and Future CMB Polarization Measurements,''
  Phys.\ Rev.\ D {\bf 84}, 043522 (2011); %  [arXiv:1106.4018 [astro-ph.CO]].
  %
 J.~Urrestilla, N.~Bevis, M.~Hindmarsh and M.~Kunz,
  %``Cosmic string parameter constraints and model analysis using small scale Cosmic Microwave Background data,''
  JCAP {\bf 1112}, 021 (2011); % [arXiv:1108.2730 [astro-ph.CO]].
  %``CMB power spectra from cosmic strings: predictions for the Planck satellite and beyond,''
  Phys.\ Rev.\ D {\bf 82}, 065004 (2010); % [arXiv:1005.2663 [astro-ph.CO]].
 %
 C.~Ringeval,
  %``Cosmic strings and their induced non-Gaussianities in the cosmic microwave background,''
 Adv.\ Astron.\  {\bf 2010}, 380507 (2010);  %   [arXiv:1005.4842 [astro-ph.CO]].
%  
 %\bibitem{Ringeval:2012tk} 
 C.~Ringeval and F.~R.~Bouchet,
  %``All sky CMB map from cosmic strings integrated Sachs-Wolfe effect,''
 Phys.\ Rev.\ D {\bf 86}, 023513 (2012). % [arXiv:1204.5041 [astro-ph.CO]].

\bibitem{Branes}
%  \bibitem{Battye:2003ks}
  R.~A.~Battye, C.~Van de Bruck and A.~Mennim,   %``Cosmological tensor perturbations in the Randall-Sundrum model: Evolution in the near brane limit,''
  Phys.\ Rev.\ D {\bf 69} 064040 (2004); % [hep-th/0308134].
%  \bibitem{Easther:2003re}
  R.~Easther, D.~Langlois, R.~Maartens and D.~Wands,   %``Evolution of gravitational waves in Randall-Sundrum cosmology,''
  JCAP {\bf 0310} 014 (2003). % [hep-th/0308078].
  %
\bibitem{Challinor}  
B.~Leong, A.~Challinor, R.~Maartens and A.~Lasenby,
  %``Brane world tensor anisotropies in the CMB,''
  Phys.\ Rev.\ D {\bf 66} 104010 (2002). %  [astro-ph/0208015].
  %%CITATION =ASTRO-PH/0208015;%%
 
 \bibitem{Zhong:2010ae} 
  Y.~Zhong, Y.~-X.~Liu and K.~Yang,   %``Tensor perturbations of $f(R)$-branes,''
  Phys.\ Lett.\ B {\bf 699}, 398 (2011) %  [arXiv:1010.3478 [hep-th]].
  
  

\bibitem{Ref5} 
A. A. Starobinsky {\it Phys. Lett.} {\bf 91B} 99 (1980);
S. Nojiri and S. Odintsov {\it Phys. Rev. } {\bf D 68} 123512 (2003); {\it Gen. Rel. Grav.}  {\bf 36} 1765 (2004);
S. M. Carroll {\it et al}, {\it Phys. Rev.} {\bf  D71} 063513 (2005);
S. Carloni, P. K. S. Dunsby, S. Capozziello and A. Troisi  {\it Class. Quant. Grav.} {\bf 22} 4839 (2005). 
J. A. R. Cembranos {\it Phys. Rev.} {\bf D 73} 064029  (2006); Phys.\ Rev.\ Lett.\  {\bf 102} 141301 (2009); 
 T. Clifton and J. D. Barrow %\The Power of General Relativity," 
{\it Phys. Rev.} {\bf D 72} 103005  (2005); %[arXiv:grqc/0509059].
  D.~S\'aez-G\'omez,   %``Modified f(R) gravity from scalar-tensor theory and inhomogeneous EoS dark energy,''
  Gen.\ Rel.\ Grav.\  {\bf 41}, 1527 (2009); %[arXiv:0809.1311 [hep-th]];
 P.~K.~S.~Dunsby {it et al.}, %E.~Elizalde, R.~Goswami, S.~Odintsov and D.S\'aez-G\'omez, %``On the LCDM Universe in f(R) gravity,''
  Phys.\ Rev.\ D {\bf 82}, 023519 (2010).   %[arXiv:1005.2205 [gr-qc]].

\bibitem{Ref6}
% S.~Tsujikawa, Lect.\ Notes Phys.\  {\bf 800}, 99 (2010). %  [arXiv:1101.0191 [gr-qc]]. 
T. Clifton {\it et al.}, %P. G. Ferreira, A. Padilla, C. Skordis, 
Phys.\ Rept.\  {\bf 513}, 1 (2012). %  [arXiv:1106.2476 [astro-ph.CO]].
% V. Faraoni %f(R)gravity, successes and challenges, ({\it Preprint} arXiv:0810.2602v1 [gr-qc]) (2008).   

\bibitem{gr-qc/0607118} 
  A.~de la Cruz-Dombriz and A.~Dobado,
  %``A f(R) gravity without cosmological constant,''
  Phys.\ Rev.\ D\ {\bf 74}, 087501  (2006)
  %[gr-qc/0607118].   %%CITATION =PHRVA,D74,087501;%%

\bibitem{Dombriz_Backreaction} A. de la Cruz-Dombriz, in preparation.

 \bibitem{Viable}
  W.~Hu and I.~Sawicki, %``Models of f(R) Cosmic Acceleration that Evade Solar-System Tests,''
  Phys.\ Rev.\ D {\bf 76}, 064004 (2007);   %[arXiv:0705.1158 [astro-ph]];
  %
%  S.~Nojiri and S.~D.~Odintsov,
  %``Modified f(R) gravity unifying R**m inflation with Lambda CDM epoch,''
  %Phys.\ Rev.\ D {\bf 77}, 026007 (2008); %[arXiv:0710.1738 [hep-th]];
  %
     L.~Pogosian and A.~Silvestri,
  %``The pattern of growth in viable f(R) cosmologies,''
  Phys.\ Rev.\ D {\bf 77}, 023503 (2008);    %[arXiv:0709.0296 [astro-ph]];
  %
S.~Capozziello and S.~Tsujikawa,   %``Solar system and equivalence principle constraints on f(R) gravity by chameleon approach,''
  Phys.\ Rev.\ D {\bf 77}, 107501 (2008). %[arXiv:0712.2268 [gr-qc]].
  
  
\bibitem{JCAP_Dombriz_2013} A. Abebe, M. Abdelwahab, A. de la Cruz-Dombriz and P. K. S. Dunsby,  Class., in preparation. 

\bibitem{SDSS} % see the webpage 
%\begin{verbatim}http://www.sdss.org/\end{verbatim}
\begin{verbatim}http://space.mit.edu/home/tegmark/sdss.html\end{verbatim}

\bibitem{Tegmark} %\cite{Tegmark:2006az}
  M.~Tegmark {\it et al.}  [SDSS Collaboration],
  %``Cosmological Constraints from the SDSS Luminous Red Galaxies,''
  Phys.\ Rev.\ D {\bf 74}, 123507 (2006). % [astro-ph/0608632].


\bibitem{Tsujikawa_BOSS} H. Okada, T. Totani, S. Tsujikawa, 	arXiv:1208.4681 [astro-ph.CO].



\bibitem{other_collab}
%\cite{Zhao:2011te}
G.~-B.~Zhao, H.~Li, E.~V.~Linder, K.~Koyama, D.~J.~Bacon and X.~Zhang,  %``Testing Einstein Gravity with Cosmic Growth and Expansion,''
  Phys.\ Rev.\ D {\bf 85}, 123546 (2012); %  [arXiv:1109.1846 [astro-ph.CO]].
 % 
  %\cite{Ross:2012sx}
  A.~J.~Ross {\it et al.},
  %``The Clustering of Galaxies in SDSS-III DR9 Baryon Oscillation Spectroscopic Survey: Constraints on Primordial Non-Gaussianity,''
  arXiv:1208.1491 [astro-ph.CO];    %%CITATION = ARXIV:1208.1491;%%
%
%\cite{Nuza:2012mw}
%\bibitem{Nuza:2012mw} 
 S.~E.~Nuza {\it et al.},
  %``The clustering of galaxies at z~0.5 in the SDSS-III Data Release 9 BOSS-CMASS sample: a test for the LCDM cosmology,''
  arXiv:1202.6057 [astro-ph.CO]; %%CITATION = ARXIV:1202.6057;%%
  %
  %\cite{Anderson:2012sa}
  % \bibitem{Anderson:2012sa} 
  L.~Anderson  {\it et al.},
  %``The clustering of galaxies in the SDSS-III Baryon Oscillation Spectroscopic Survey: Baryon Acoustic Oscillations in the Data Release 9 Spectroscopic Galaxy Sample,''
  arXiv:1203.6594 [astro-ph.CO];   %%CITATION = ARXIV:1203.6594;%%
  %
  %\cite{Ross:2012qm}
  %\bibitem{Ross:2012qm} 
  A.~J.~Ross {\it et al.},
  %``The clustering of galaxies in the SDSS-III Baryon Oscillation Spectroscopic Survey: Analysis of potential systematics,''
  arXiv:1203.6499 [astro-ph.CO]. %%CITATION = ARXIV:1203.6499;%%
  %
  %
  %\cite{Scoccola:2012ny}
  %\bibitem{Scoccola:2012ny} 
  C.~G.~Scoccola {\it et al.},
  %``The clustering of galaxies in the SDSS-III Baryon Oscillation Spectroscopic Survey: constraints on the time variation of fundamental constants from the large-scale two-point correlation function,''
  arXiv:1209.1394 [astro-ph.CO]. %%CITATION = ARXIV:1209.1394;%%
  


%%%%%%%%%%%%%%%%%%%%%%%%%
%\bibitem{CAMB}  
\bibitem{CAMB_Lewis} A. Lewis, A. Challinor, and A. Lasenby, Astrophys. J. 538, 473 (2000);  A. Lewis, PhD Thesis, University of Cambridge (2000).

% \bibitem{CMB-Easy} CMB Easy reference.

\bibitem{CMB-Fast} 
    U. Seljak and M. Zaldarriaga, 1996 ApJ, 469, 437; 
    M. Zaldarriaga, U. Seljak and E. Bertschinger, 1998 ApJ, 494, 491; 
    M. Zaldarriaga and U. Seljak, 2000 ApJS, 129, pp 431-434.
%${\rm http://lambda.gsfc.nasa.gov/toolbox/tb\_cmbfast\_ov.cfm}$
    
 \bibitem{WMAP}   N.~Jarosik, C.~L.~Bennett, J.~Dunkley, B.~Gold, M.~R.~Greason, M.~Halpern, R.~S.~Hill and G.~Hinshaw {\it et al.},
  Astrophys.\ J.\ Suppl.\  {\bf 192}, 14 (2011);  %[arXiv:1001.4744 [astro-ph.CO]].
  E.~Komatsu {\it et al.}  [WMAP Collaboration], Astrophys.\ J.\ Suppl.\  {\bf 192}, 18 (2011). %[arXiv:1001.4538 [astro-ph.CO]].
  
\bibitem{Planck} Planck Collaboration, ESA-SCI-2005-1, arXiv: astro-ph/0604069.



  %%%%%%%%%%%%%%  %%%%%%%%%%%%%%  %%%%%%%%%%%%%  %%%%%%%%%%%%%%  %%%%%%%%%%%%%%   
%
\end{thebibliography}

\documentclass[graybox]{svmult}

\usepackage{mathptmx}       
\usepackage{helvet}         
\usepackage{courier}        
\usepackage{type1cm}        
\usepackage{makeidx}         
\usepackage{graphicx}        
\usepackage{multicol}        
\usepackage[bottom]{footmisc}

\makeindex             

\begin{document}

\title*{Towards degeneracy problem breaking by large scale structures methods}
\author{\'Alvaro de la Cruz Dombriz}
\institute{Astronomy, Cosmology and Gravity Center (ACGC) and
Department of Mathematics - Applied Mathematics,
University of Cape Town, Rondebosch 7700,  South Africa. \\ \email{alvaro.delacruzdombriz@uct.ac.za} 
}
%
%
\maketitle

\abstract{An arguable aspect of the modified gravity theories is that many of them present the so-called {\it degeneracy problem}. For instance, the cosmological evolution, gravitational collapse and the main features of standard black-hole configurations, can be mimicked by many of those theories. In this communication we revise briefly the appropriate observable quantities to be measured in order to discard alternative theories to $\Lambda$CDM, such as the observed growth of scalar perturbations with Sloan data and the CMB tensor perturbations evolution.
}

\section{Introduction}
\label{sec:1}

Modified gravity \cite{review} 
has been shown to be able to mimic both the dark energy (DE) and the inflationary eras \cite{unification}.  However the use of large scale observations, such as Ia type supernova, baryon acoustic oscillations, or the cosmic microwave background (CMB), which only depend upon the expansion history of the Universe is not enough to determine uniquely the nature and the origin of DE. Let us rephrase the argument: 
identical cosmological background evolutions 
can be explained by a pleiad of theories. This is the so-called {\it degeneracy problem}, whose breaking requires measurements 
not only sensitive to the 
cosmological expansion but, for instance, the evolution of scalar perturbations \cite{Perturbations}, the stability of cosmological solutions when subjected to small perturbations \cite{Dombriz-Saez} and the existence of General Relativity (GR)-predicted astrophysical objects such as black holes \cite{BH}. Finally, the study of CMB tensor perturbations may also shed some light about the viability of modified gravity theories \cite{CMB-ERE2011, PRD-CMB-Dombriz, Strings, Branes, Challinor, Zhong:2010ae}.  

In this realm, the simplest and in fact the most studied modification of  the Hilbert-Einstein action is generalized to a general function of the Ricci scalar $R$, dubbed $f(R)$ gravity theories  \cite{Ref5}-\cite{Ref6} whose action can be written as 
>%
 \begin{eqnarray}
 {\cal{A}}=\frac{1}{16\pi G}\int{{\rm d}^{4}x\sqrt{-g}\left(R+f(R)+2{\cal{L}}_{m}  \right) }\;,
 \end{eqnarray}
 where the symbols hold their usual meanings.
In addition to reproducing the entire cosmological history \cite{gr-qc/0607118} and despite some shortcomings  \cite{Ref6}, these theories may behave quite well on local scales, where the GR limit must be recovered \cite{Viable}. As for any alternative theory of gravity, 
in $f(R)$ theories, the density contrast evolution, the CMB perturbations and the backreaction mechanism \cite{Dombriz_Backreaction}, if the latter is assumed to be true, need to be studied 
in order to unveil the potential distinct features of these scenarios. 
In the present investigation we sketch the main 
features and steps to study the two first issues in $f(R)$ theories. 

\section{Scalar perturbations in $f(R)$ theories}
\label{sec:2}
The density contrast evolution for $f(R)$ theories obeys a fourth-order differential equation \cite{PRD_Dombriz_2008}.
The resulting equation for the density contrast $\delta$  can be written as follows:
\begin{eqnarray}
&&\beta_{4,f}\delta^{iv}+\beta_{3,f}\delta'''
+(\alpha_{2,EH}+\beta_{2,f})\delta''
+(\alpha_{1,EH}+\beta_{1,f})\delta'+ 
(\alpha_{0,EH}+\beta_{0,f})\delta \,=\,0\nonumber\\
&&
\label{density_contrast_eqn}
\end{eqnarray}
where the coefficients $\beta_{i,f}$  $(i\,=\,1,...,4)$ 
involve terms that disappear for $f(R)$ functions linear in $R$ (i.e., GR) whereas $\alpha_{i,EH}$  $(i\,=\,0,1,2)$ involve the linear part in $R$ of $f(R)$.
Thus, the quasi-static limit ($k>>\mathcal{H}$) of (\ref{density_contrast_eqn}) becomes  \cite{PRD_Dombriz_2008}
\begin{eqnarray}
&&\delta''+\mathcal{H}\delta'+\frac{(1+f_{R})^{5} \mathcal{H}^{2} 
(-1+\kappa_1)(2\kappa_1-\kappa_2)-\frac{16}{a^8}
 f_{RR}^{4}(\kappa_2-2)k^{8}8\pi G \rho_{0}a^2}
{(1+f_{R})^{5}(-1+\kappa_1)+\frac{24}{a^8}f_{RR}^{4}(1+f_{R})
(\kappa_2-2)k^{8}}\delta \,=\,0\nonumber\\
&&
\label{density_contrast_eqn_QSA}
\end{eqnarray}
Contrarily to its counterpart for $\Lambda$CDM, the 
coefficients in (\ref{density_contrast_eqn}) depend both upon the model under consideration and the wavenumber $k$. This fact gives rise to
$k$-dependent transfer functions that may alter dramatically the matter power spectra \cite{PRD_Dombriz_2008, PRL_Dombriz_2009, JCAP_Dombriz_2013}.
Available  data \cite{SDSS} using luminous red galaxies in the Sloan Digital Sky Survey (SDSS) were able to measure the large-scale real-space power spectrum. 
These measurements were used to sharpen the constraints on cosmological parameters \cite{Tegmark} and may be straightforwardly compared with the predictions made by gravity theories
\cite{PRL_Dombriz_2009, Tsujikawa_BOSS}. Very recently a full study for the $R^n$ models \cite{JCAP_Dombriz_2013} have stressed the importance of the initial conditions in the perturbed equations
which determine the evolution of the transfer function. Consequently, this method provides an excellent arena to impose 
tight constraints for modified gravity models that are claimed to be valid 
once compared with existing and future data \cite{other_collab}.

\section{CMB perturbations in $f(R)$ theories}
\label{sec:2}
The study of the CMB tensor perturbations in modified gravity theories has not received much interest in comparison 
with the scalar counterpart. This fact has laid in the difficulty of obtaining 
the required tensor perturbed equations which are in general of higher order. An alternative route in order to circumvent this difficulty consists of tackling the problem 
by using the simulations performed by several codes available such as CAMB \cite{CAMB_Lewis} 
based upon modifications of CMBFast \cite{CMB-Fast}.

Different attempts were made for several modified gravity scenarios \cite{Strings} 
but  most of the attention was devoted to the study of the tensor perturbations evolution in the brane-world theories context 
\cite{Branes, Challinor}. 
Finally, with regard to $f(R)$ fourth order gravity theories, the only attempts to encapsulate 
 the main features of tensor perturbation were made in \cite{Zhong:2010ae} and more recently in \cite{CMB-ERE2011, PRD-CMB-Dombriz}. 
The authors of the first investigation analyzed the tensor perturbations of 
flat thick domain wall branes in  $f(R)$ gravity. They showed that under the transverse and traceless gauge, the metric perturbations decouple from the perturbation of the background scalar field which generates the brane. Authors in \cite{CMB-ERE2011, PRD-CMB-Dombriz} addressed for the first time in literature the tensor perturbations full calculations for the $f(R)$ gravity theories in the metric formalism and Jordan frame. These general results were applied to $R^n$ models for different values of $n$ 
%
describing the features that may distinguish those models from Concordance model predictions. This implementation proved the importance of considering the correct background when 
alternative theories of gravity are subjected to this kind of analyses since a relevant contribution to the  $c_{l}^{\rm TT}$ and  $c_{l}^{\rm EE}$ CMB coefficients 
comes from the background implementation.

Thus, exclusions tests for $f(R)$ models can be performed since data for $c_{l}^{\rm TT}$ are already available from WMAP \cite{WMAP} once the scalar contribution are also included. With respect to $c_{l}^{\rm EE}$  once Planck \cite{Planck} measurements are ready, some data may be compared with theoretical predictions.


%
\begin{acknowledgement}
The author acknowledges financial support from URC and NRF (South Africa), MICINN (Spain) project numbers FIS2011-23000, FPA2011-27853-C02-01 and Consolider - Ingenio MULTIDARK CSD2009-00064. 
\end{acknowledgement}
%

%
%
%

%
%

\end{document}